### v2

# Imaging-spectroscopy diagnosis of the giant sloshing spiral in the Virgo cluster with the Einstein Probe Follow-up X-ray Telescope

11institutetext: State Key Laboratory of Particle Astrophysics, Institute of High Energy Physics, Chinese Academy of Sciences, Beijing 100049, China 22institutetext: School of Physics and Astronomy, Beijing Normal University, Beijing 100875, China 33institutetext: Department of Astronomy, Tsinghua University, Beijing 100084, China 44institutetext: Max Planck Institute for Extraterrestrial Physics, Giessenbachstrasse 1, 85748 Garching, Germany 55institutetext: Institute for Frontiers in Astronomy and Astrophysics, Beijing Normal University, Beijing 102206, China 66institutetext: National Astronomical Observatories, Chinese Academy of Sciences, 20A Datun Road, Beijing 100101, People’s Republic of China 77institutetext: Department of Astronomy, Xiamen University, Xiamen 361005, China 88institutetext: School of Astronomy and Space Science, University of Chinese Academy of Sciences, 19A Yuquan Road, Beijing 100049, People’s Republic of China 99institutetext: Computing Center, Institute of High Energy Physics, Chinese Academy of Sciences, Beijing 100049, China
Imaging-spectroscopy diagnosis of the giant sloshing spiral in the Virgo cluster with the Einstein Probe Follow-up X-ray Telescope
X. Zheng 1122    S. Jia 11    C. Li 11    Y. Chen 11    H. Yu 22    H. Feng 11    D. Xu 33    A. Liu 224455    L. Song 11    C. Liu 11    F. Lu 11    S. Zhang 11    W. Yuan 66    J. Sanders 44    J. Wang 77    T. Chen 11    C. Cui 66    W. Cui 11    W. Feng 11    N. Gao 11    J. Guan 11    D. Han 11    D. Hou 11    H. Hu 66    M. Huang 66    J. Huo 11    C. Jin 556688    M. Li 11    W. Li 11    Y. Liu 66    L. Luo 11    J. Ma 11    G. Ou 99    H. Pan 66    H. Wang 11    J. Wang 11    J. Wang 11    Y. Wang 11    J. Xu 11    Y. Xu 66    X. Yang 11    Y. Yang 11    H. Zhang 99    J. Zhang 11    M. Zhang 66    Z. Zhang 11    H. Zhao 11    X. Zhao 11    Z. Zhao 11    P. Zhu 11    Y. Zhu 11
(Received September 30, 20XX)
We performed deep X-ray observations of the Virgo cluster using the Einstein Probe Follow-up X-ray Telescope (EP-FXT) with a total exposure of 295 ks. Taking advantage of the large field of view (FoV) and low particle background of EP-FXT, the image reveals a giant spiral feature connecting the cold fronts in the northwest and southeast, forming a coherent structure consistent with previous XMM-Newton and Suzaku findings. Furthermore, we present two-dimensional maps of temperature, metallicity, and entropy in the Virgo Cluster, covering a FoV of approximately $28.5^{\prime}$. These maps clearly depict a spiral structure characterized by high density, low temperature, high metallicity, and low entropy. These results support the scenario in which the spiral morphology originates from gas sloshing induced by a minor merger. In addition, the temperatures measured with EP-FXT show reasonable consistency with those obtained from XMM-Newton, within uncertainties.
Key Words.:
EP-FXT – X-ray – galaxy cluster – Virgo Cluster
1 Introduction
Galaxy clusters, composed of hundreds to thousands of galaxies bound by gravity, are the largest building blocks of the cosmic web (Kravtsov & Borgani 2012; Vikhlinin et al. 2014). The X-ray radiation from clusters predominantly arises from the intracluster medium (ICM), a hot, diffuse plasma where a wealth of complex structures can be seen, such as shocks, bubbles, and cold fronts, which trace the past and ongoing violent dynamic processes in the system (Markevitch & Vikhlinin 2007). Cold fronts, distinguished from shock fronts, are sharp interfaces separating brighter, denser, and cooler gases from the other side with a continuous pressure across the interface (Owers et al. 2009; de Plaa et al. 2010; Ettori et al. 2013; Ghizzardi et al. 2010).
Cold fronts or even spectacular spiral structures have been widely identified in nearby clusters (Keshet 2012). It is proposed that the cold front originates from a displacement of cold gas at the cluster core, triggered by gravitational perturbations—such as those induced by minor mergers (Ascasibar & Markevitch 2006; ZuHone et al. 2010; Rossetti 2013; Markevitch et al. 2001; Roediger et al. 2011; Markevitch & Vikhlinin 2007). In relaxed galaxy clusters, minor mergers disturb the central gas, initiating sloshing motions that produce multiple cold fronts at various locations over time (Roediger et al. 2011; Liu et al. 2018). These sloshing motions can evolve into intricate spiral structures, visible in X-ray images as concentric edges (Ascasibar & Markevitch 2006; Sonkamble et al. 2024), such as the case of A795 (Kadam et al. 2024a). Alternative explanations for the sloshing-induced spiral include two-phase gas flows with differing velocities (Keshet 2012) and the rotational motion of an in-falling gas flow around a moving brightest cluster galaxy (BCG) due to the Coriolis force (Inoue 2022).
Despite the detection of a variety of cold fronts, those manifesting themselves as a clear and complete spiral structure are rare, mainly due to difficulties in revealing large-scale and low-contrast structures with current X-ray telescopes. So far, remarkable examples include those seen in A496, A795, A1775, A2029, A2052, A2566, A2657, Centaurus, NGC 7618/UGC 12491, Perseus, and Virgo (Walker et al. 2017; Fabian et al. 2006; Roediger et al. 2012; Ghizzardi et al. 2014; Kadam et al. 2024a; Botteon et al. 2024; Machacek et al. 2023; Blanton et al. 2011; Paterno-Mahler et al. 2013; Sanders et al. 2016b; Botteon et al. 2021; Kadam et al. 2024b), mainly thanks to mosaic observations with XMM-Newton and Chandra. Detection and detailed investigations of large-scale sloshing spirals are essential for advancing our understanding of the physical mechanism of sloshing and dynamical evolution of galaxy clusters.
Because of its proximity, the Virgo cluster serves as an excellent target for in-depth studies on this topic. Within the Virgo cluster, M87 stands out as the largest and brightest elliptical galaxy, with a redshift of 0.0042. It has been extensively studied across multiple wavelengths, providing valuable insights into the interaction between its central supermassive black hole ($3.2\times 10^{9}M_{\odot}$ within 3.5 pc of its center) and the surrounding ICM (Harms et al. 1994; Simionescu et al. 2010; Bahcall & Sarazin 1977). Early X-ray observations with ROSAT, XMM-Newton, and Chandra have provided definitive evidence for active galactic nucleus (AGN) feedback surrounding the central galaxy M87 (Böhringer et al. 2001; Molendi & Gastaldello 2001; Bohringer et al. 1995; Forman et al. 2005; Sanders et al. 2016a). In addition, several distinct structures within M87 have been discovered, such as two X-ray arms, bubbles, and cavities (Sanders et al. 2016a). X-ray observations have revealed multiple sloshing cold fronts in the Virgo Cluster. XMM-Newton identified cold fronts at 33 kpc and 90 kpc, and revealed a spiral structure extending to approximately 200 kpc (Simionescu et al. 2010; Roediger et al. 2011). Suzaku later detected additional outer cold fronts at 233 and 280 kpc and carried out detailed radial analyses of their thermodynamic properties, further confirming that gas sloshing extends well beyond the core of the cluster (Simionescu et al. 2017). Although previous studies have mapped the two-dimensional thermodynamic properties of the ICM in the central region of M87 Gatuzz et al. (2022), the temperature, entropy, and metallicity distributions in the outer spiral arm remain poorly characterized. While eROSITA is capable of covering the extended spiral structure on the outskirts of Virgo (McCall et al. 2024), its relatively short exposure time has limited detailed analysis of the associated physical properties. To more comprehensively investigate the role of gas sloshing and cold fronts in the outskirts of galaxy clusters, it is important to conduct observations with an X-ray telescope that offers a larger FoV and sufficient sensitivity.
Benefiting from its large FoV and low particle background, the Follow-up X-ray Telescope onboard the Einstein Probe (EP-FXT)(Yuan et al. 2022, 2025) is well suited for detecting large-scale, low surface brightness extended structures. In this paper, we present deep EP-FXT observations of the Virgo cluster. Our data support the detection of the giant sloshing spiral and enable spatially resolved analyses of its thermodynamic properties. The paper is organized as follows. We describe the observations in Sect. 2, with imaging results presented in Sect. 3 and spatially-resolved spectral results in Sect. 4. We discuss the physical implications in Sect. 5 and summarize our findings in Sect. 6.
2 Observations and data reduction
EP-FXT is one of the two payloads onboard the EP mission,which is an X-ray telescope as a result of China-Europe collaboration, launched on January 9, 2024. EP-FXT is capable of conducting detailed observations of targets triggered by the Wide-field X-ray Telescope (EP-WXT), as well as preselected targets designated from the ground. EP-FXT comprises two identical modules, FXT-A and FXT-B, each equipped with Wolter-I type X-ray optics featuring a focal length of 1.6 m and 54 nested gold-coated nickel mirror shells. These modules are paired with PNCCD detectors as focal-plane instruments, enabling precise measurements of photon energy, position, and timing (Cui et al. 2023). The device features a pixel array of 384 $\times$ 384, with an angular resolution of approximately $9.6^{\prime\prime}$ per pixel. EP-FXT covers an energy range of 0.3–10 keV, with a FoV of $1^{\circ}\times 1^{\circ}$, a spatial resolution of $22^{\prime\prime}$ in terms of half-power diameter (HPD), and an effective area of approximately 300 $cm^{2}$ at 1 keV (one unit). In addition, the particle background of EP-FXT is about one-fifth that of eROSITA, and its low background characteristics make it particularly well-suited for observing low-brightness diffuse emissions (Zhang et al. 2022).
As the first-light object and a calibration source for EP-FXT, the Virgo cluster and its central galaxy M87 have been extensively observed. Due to the large FoV of EP-FXT, a single observation can effectively cover the outskirts of M87. All observations were performed in full-frame (FF) mode, using either the thin or medium filter. Three of the FXT-B observations were conducted in a closed mode, rendering the corresponding data unusable. As a result, only FXT-A data were used, resulting in a total effective exposure time of 295 ks. (see an observational log in Table LABEL:label1).
Figure 1: Combined EP-FXT X-ray image of the Virgo cluster in the 0.3-2.5 keV band. The image has been vignetting corrected, background and point-source subtracted, flattened, and adaptively smoothed. The white circle marks the background region. The bar has a length of 10. The arrows point to north and west.
We used the latest version 1.10 of the Follow-up X-ray Telescope Data Analysis Software (FXTDAS) [^1] http://epfxt.ihep.ac.cn/analysis developed at the EP-FXT Science Center for data reduction (Zhao et al. 2025). New event files were generated through the fxtchain pipeline, and the out-of-time events were identified using fxtootest, which were then removed from the data with ftimgcalc [^2] https://heasarc.gsfc.nasa.gov/docs/software/lheasoft/help/heatools.html. When high-energy cosmic rays pass through space telescopes, they release energy onto the detectors, exciting fluorescence lines that are not modulated by the vignetting effect. To correct for this background contribution, fxtbkggen was used to generate particle backgrounds for each observation. We removed the particle background from each image and then used the reproject_image [^3] Derived from the Chandra Interactive Analysis of Observations (CIAO) version 4.16.0, available at https://cxc.cfa.harvard.edu/ciao/ tool to merge all exposure images along with their corresponding particle-background-subtracted images. Subsequently, we used the merged exposure maps to correct for the vignetting effect in the particle background-subtracted X-ray images. Following this, the data from ObsID 13600005125, with the particle background already subtracted, were selected as the reference for the diffuse background. Given that the primary focus of this study is M87, we selected a background region in close proximity to the M87. This approach reduces the potential to underestimate the diffuse background component surrounding M87, thereby facilitating more precise background subtraction and strengthening the overall robustness of the analysis. We combined multiple observations, subtracted the diffuse background, and normalized the image by the maximum exposure time to produce a calibrated X-ray image. We used wavedetect3 with scales = ”4 8 16 32” and the default sigthresh to detect and remove point sources. The voids were filled using dmfilth3 with the POISSON method and a random seed of 0 to ensure reproducibility. Adaptive smoothing was then applied using dmimgadapt with Gaussian kernels, smoothing scales from 0.1 to 10 pixels, and a minimum of 30 counts per kernel. The final X-ray image in the energy range of 0.3-2.5 keV shown in Fig. 1.
Furthermore, to compare with previous results, we reanalyzed a Chandra observation (ObsID 5826) with an exposure of 126 ks, and an XMM-Newton observation (ObsID 0803670601, EPIC-MOS2) with an exposure of 65 ks. The data were reprocessed using the XMM-Newton Science Analysis System (SAS  [^4] https://www.cosmos.esa.int/web/xmm-newton/sas), and time intervals affected by flares were removed through high-energy light curve screening. In the subsequent XMM-Newton analysis, blank-sky background files were employed to improve the accuracy of background estimation and spectral analysis.
Table 1: EP-FXT observations of the Virgo cluster.
ObsID RA DEC Date Exposure time (ks) filter Instrument
08500000007 187.71 12.39 2024-02-22 3.38 thin A
08500000008 187.71 12.39 2024-02-22 20.46 medium A
13600005124 187.71 12.39 2024-03-17 77.73 thin A/B
13600005125 (offset) 187.01 12.87 2024-03-18 23.78 thin A/B
13600005132 187.71 12.39 2024-04-05 28.08 medium A
13600005431 187.71 12.39 2024-04-16 18.15 medium A/B
13600006200 187.71 12.39 2024-05-20 97.17 medium A/B
11904194486 187.71 12.39 2024-06-15 50.18 thin & medium A/B
3 Imaging analysis
To visualize faint structures in the image, we applied two different techniques for image enhancement.
Figure 2: Adaptive GGM filtered image in the central Virgo cluster in 0.3-2.5 keV with SNR = 70. Prominent features are labeled. The two dashed line to mark the giant spiral are identical to those in Fig. 4.
Gradient filtering enhances the visibility of both sharp and flat image features, thereby facilitating their detection and analysis (Sanders et al. 2016b, a, 2022). Because of the high noise in X-ray image gradient measurements within low-brightness regions, the adaptive Gaussian Gradient Magnitude (GGM) filtering is employed for dynamic scale selection. This method dynamically adjusts filter parameters based on local gradients, enabling better capture of image textures and edges (Sanders et al. 2022). Additionally, adjusting the signal-to-noise ratio (SNR) helps smooth the image, enhances the SNR, and reveals finer details. Using these techniques, we applied the adaptive GGM filter to the vignetting-corrected and background-subtracted image in the 0.3–2.5 keV energy range, see Fig. 2.
The image clearly shows two X-ray arms, with the southwestern arm exhibiting a clockwise rotation toward the southeast (Forman et al. 2007). This rotational feature may be attributed to the interaction between relativistic jets from the AGN and the ICM (Forman et al. 2005). Furthermore, the image also identifies a cavity, a shock to the northwest, and cold fronts to the southeast and northwest. These features are consistent with the analysis by Sanders et al. (2016a) using the Chandra data. We also identified a region of significantly reduced surface brightness outward from the southwestern X-ray arm, which may indicate the presence of a cavity. The outskirts of M87 exhibit a prominent spiral morphology, extending from the northwest to the southeast, with similar spiral features also visible in the XMM-Newton mosaic image (Roediger et al. 2011; Simionescu et al. 2010). We then generate a residual image to better visualize this large-scale faint feature.
First, we need to estimate the first-order radial profile using the well-known double $\beta$ model and subtract it. As shown in Fig. 3, the red solid line represents the best-fit double $\beta$ model to the EP-FXT data (black points). The blue points represent the XMM-Newton data and the green points represent the Chandra data, for comparison. As shown, apart from the central region affected by the AGN, the EP-FXT data are in agreement with the double $\beta$ model and exhibit consistency with the radial profiles measured by the three telescopes. Furthermore, thanks to EP-FXT’s large FoV, its observational range extends to regions up to $30^{\prime}$ from the outskirts of M87.
Figure 3: Normalized surface brightness profiles measured with EP-FXT, XMM-Newton, and Chandra. The EP-FXT data are fitted with a double $\beta$ model, with individual components displayed in dashed yellow curves.
Figure 4: Residual image obtained by subtracting the image in Fig. 1 with the elliptical $\beta$ model. A Gaussian smooth with a kernel size of 7 pix is applied (1 pixel = $9.6^{\prime\prime}$). Four regions of interest (NW, SE, SW and NE) are marked. The white dashed lines mark the position of radial discontinuity. The blue dashed lines mark the positions of the giant spiral arm.
To characterize the radial profile, we used Sherpa (version 4.16.0 within CIAO) [^5] https://cxc.cfa.harvard.edu/sherpa/ to model the surface brightness. To account for the apparent elongated morphology as seen from the image, we fitted an elliptical $\beta$ model to the data; the best‑fit parameters are listed in Table 2. We then subtracted this model from the original image to obtain the residual image shown in Fig. 4.
Table 2: Best‐fit parameters of the elliptical $\beta$ model from Sherpa.
$r_{0}$ (arcsec) $\mathrm{ellip}\leavevmode\nobreak\ (10^{-2})$ $\theta$ (deg) $A$ (counts pixel-1) $\alpha\leavevmode\nobreak\ (10^{-2})$
$90.58^{+0.39}_{-0.39}$ $10^{+0.04}_{-0.04}$ $248.13^{+0.12}_{-0.12}$ $4528.09^{+17.9}_{-17.9}$ $74^{+0.09}_{-0.09}$
The residual image reveals a few patterns, including X-ray arms and discontinuities which indicate cold fronts or shocks. Notably, we identified a ring-like structure outside the X-ray arms, which displays a break in the southern part. This may be attributed to interactions between the AGN jets and the ICM (Forman et al. 2005). Most importantly, the large-scale spiral structure with excessive brightness is evident on the outskirts, spiraling from the cold front in the northwest to the southern cold gas clump, marked with blue dashed lines in Fig. 4 and white dashed lines in Fig. 2. Consistent with the mosaic XMM-Newton results (Simionescu et al. 2010), the EP-FXT observations clearly reveal the spiral features as continuous and well-defined across the entire FoV, likely owing to the instrument’s low particle background and wide spatial coverage.
The northwestern outer edge (a cold front) of the spiral lies approximately $18.3^{\prime}$ from the center. The structure rotates counterclockwise toward the south, with its southern inner edge located approximately $17.5^{\prime}$ from the center.
Figure 5: An example of the EP-FXT spectrum from a contour-binning region in M87. The black and red points represent the spectral data from FXT-A and FXT-B for ObsID 13600005124, with the corresponding background from ObsID 13600005125 subtracted, and the solid line represents the best-fit model. The green and blue points represent the spectra from FXT-A and FXT-B for ObsID 13600005125, respectively. The light cyan and orange represent the particle backgrounds for ObsIDs 13600005124 and 13600005125, respectively.
Figure 6: Temperature, pressure, entropy, and metallicty maps measured with EP-FXT, pixelated with the contour-binning method. The white dashed curves mark the spiral structure as in Fig. 4.
4 Imaging-spectral analysis
The X-ray image obtained with EP-FXT was divided into 142 regions using the contour-binning software (Sanders 2006) with a SNR of 300. We adopted a flat-weighting assumption, assuming that the X-ray emission within each region is spatially uniform, for the Ancillary Response Files (ARFs) calculation. The ARFs were generated using the EP-FXT response tool fxtarfgen. During the calculation, the parameter extend was set to 1 to account for the extended nature of the source, thus more accurately capturing the spatial distribution of the emission within each extraction region. The spectra were grouped using grppha with a minimum of 10 counts per bin to ensure the applicability of the C-statistic. To minimize the impact of background contamination on the source spectra and improve the accuracy of the spectral analysis, we selected observations of the M87 source and its corresponding background, ensuring similar observation times, the same filter settings, and longer exposure durations; specifically, ObsID 13600005124 and ObsID 13600005125. In the spectral analysis, both the particle background and the diffuse background were taken into account. We calculated the particle background for both observations, with a ratio of $1.05\pm 0.03$, indicating that the particle background levels are nearly identical. Therefore, we selected the spectra from the same CCD region in ObsID 13600005125 to serve as the background for ObsID 13600005124. For regions in ObsID 13600005125 that are contaminated by bright point sources or extended sources, background data from nearby uncontaminated areas were used instead. Given that our analysis focuses on M87, we intentionally retained the Virgo cluster component in the background data because it more accurately reflects the diffuse emission environment surrounding the galaxy.
Spectral fitting of the combined FXT-A and FXT-B spectra in each region was performed using XSPEC version 12.14.1 over the 0.5–7.0 keV energy range. The diffuse ICM thermal emission was described with the APEC model (version 3.0.9), and the Galactic hydrogen absorption was accounted for using the TBABS model, adopting the Solar abundance table from Anders & Grevesse (1989), with the redshift fixed at 0.0042 and the hydrogen column density fixed at $2.11\times 10^{20}\mathrm{cm}^{-2}$ (Willingale et al. 2013). Other parameters are set to be free. We examined the C-Stat/d.o.f. values for all regions, which were consistently around 1.2. We plotted the spectrum of one region, as shown in Fig. 5. The region belongs to the contour-binning partition and is located in a region approximately $6.6^{\prime}$ from the cluster center. The red and black points represent the spectral data from FXT‐A and FXT‐B in the source region (with the diffuse background subtracted), while the solid line denotes the best‐fit model. The C-Stat/d.o.f. for FXT-A and FXT-B are 246.74/287 and 295.70/299, respectively. The green and blue points display the spectral data from FXT‐A and FXT‐B in the diffuse background region, which is co-located on the CCD with the source region. In addition, the light cyan and orange indicate the particle backgrounds for the source and diffuse background regions, respectively.
4.1 The map of physical properties
The temperature map (see Fig. 6) reveals a central region in M87 with temperatures below 2.5 keV, along with two distinct low-temperature regions in the northwest and southeast. Sharp temperature increases are seen exterior to these regions, coincident with the locations of two concentric, arc-shaped cold fronts. A low-temperature belt extends outwards from the central region, connecting the low-temperature regions in the southeast and northwest, and spiraling counterclockwise from the northwest, forming a large-scale low-temperature spiral structure. In addition, a region with a lower temperature, compared to the surrounding areas, is located on the outer side of the southwest arm. A detailed analysis is presented in Sect. 4.2.
The metallicity are obtained from the APEC model. The metallicity is lower within the two X-ray arms but higher to the north and south of the center, which is consistent with the findings in Simionescu et al. (2007); Gatuzz et al. (2022). Additionally, beyond $12^{\prime}$, the metallicity in the northwest is significantly higher than that in the southeast, likely due to the accumulation of cold gas in the northwest. In addition, the metallicity of the spiral structure displays higher values, further supporting the existence of the spiral structure.
To investigate its thermodynamic properties, we examine the distributions of pseudo-pressure and pseudo-entropy. The thermal pressure and entropy can be defined (Sasaki et al. 2016) as $P=kTn_{\textnormal{e}}$ and $S=kTn_{\textnormal{e}}^{-2/3}$, respectively, where $n_{e}$ is the electron density and $kT$ is the temperature. Based on the emission measure obtained from spectral fits, we obtained pseudo-entropy and pseudo-pressure. we convert $P$ and $S$ to $P\propto kT{\rm norm}^{0.5}$ and $S\propto kT{\rm norm}^{-1/3}$, respectively (Fabian et al. 2006). Pseudo-pressure and pseudo-entropy maps are also displayed in Fig. 6. In the central region of M87, the pressure map reveals decreasing pressure along the eastern and southwestern X-ray arms, and increasing pressure in the northwest and southeast. The entropy map shows low-entropy features coincident with both X-ray arms, exhibiting an entropy asymmetry towards the northwest and southeast near the core. These results are consistent with those reported in Simionescu et al. (2007). The outer spiral structure also displays lower entropy values, further supporting the existence of the spiral structure.
To sum up, the EP-FXT observations revealed the spiral structure with high completeness and clarity. Moreover, we systematically mapped its thermodynamic properties—including temperature, metal abundance, and entropy—through two-dimensional distributions. These results provide detailed and spatially resolved evidence for the existence and evolution of the spiral structure.
4.2 Sectorial radial profiles
In the residual map (see Fig. 4), we defined four sectorial regions of interest (NW, SE, NE and SW). Using Fig. 1, radial profiles of the surface brightness and temperture are extracted in each region (see Fig. 7). The surface brightness profiles near the cold fronts are fitted with a broken power-law. The best‐fit parameters of the broken power‐law model are listed in Table 3.
In the northeast, two surface brightness drops are detected at $R=7.44\pm 0.04^{\prime}$ and $R=21.40\pm 0.14^{\prime}$, respectively. At $R=21.40\pm 0.14^{\prime}$, a modest increase in temperature is found from $2.16^{+0.11}_{-0.10}\leavevmode\nobreak\ \mathrm{keV}$ to $2.50^{+0.19}_{-0.16}\leavevmode\nobreak\ \mathrm{keV}$ at the $1.7\sigma$ level (The uncertainties are quoted at the $1\sigma$ level). While the temperature change alone is not highly significant, the corresponding surface brightness discontinuity and the location in the outskirts of M87—where the SNR is lower—suggest that this feature likely marks the outer boundary of the northeastern spiral structure.
In the northwest, a drop in surface brightness is found at $R=18.33\pm 0.08^{\prime}$, accompanied by a rise in temperature from $2.39^{+0.06}_{-0.07}$ to $2.70^{+0.12}_{-0.14}\leavevmode\nobreak\ \mathrm{keV}$, corresponding to a significance of approximately $2.2\sigma$. Despite the moderate significance of the temperature jump, the presence of a surface brightness edge, coupled with the low SNR in the outskirts, supports the classification of this feature as a cold front, in agreement with previous findings.
In the southeast, a cold front is identified at $R=6.76\pm 0.03^{\prime}$, indicated by a drop in surface brightness and a temperature increase from $2.43^{+0.03}_{-0.02}$ to $2.58^{+0.04}_{-0.03}\leavevmode\nobreak\ \mathrm{keV}$ at the $3.3\sigma$ level. At $R=17.52\pm 0.22^{\prime}$, the surface brightness shows a mild discontinuity, transitioning from a deficit to an excess in the residual map (see Fig. 4). This feature may correspond to the inner boundary of the southeastern portion of the spiral structure.
A region of excess surface brightness is detected in the southwest at $R=9.75\pm 0.07^{\prime}$, potentially associated with an arc-shaped feature from past AGN activity (Forman et al. 2005). The surface brightness discontinuity is described by a broken power-law model, and a corresponding radial temperature increase further supports the possible presence of a weak cold front.
We compared the radial temperature profile derived from EP-FXT observations with the results reported by Gatuzz et al. (2022) based on XMM-Newton data. The two profiles show a broadly consistent trend, with comparable shapes and temperatures ranging from approximately 2 to 3 keV.
Table 3: Best-fit broken power-law model parameters.
region $R_{\rm break}$ (arcmin) $I_{0}$ ( $10^{-2}$ cts s-1 arcmin-2) $\alpha_{1}$ $\alpha_{2}$
NW $18.33\pm 0.08$ $1.29\pm 0.02$ $2.27\pm 0.08$ $5.85\pm 0.23$
SE $6.76\pm 0.03$ $2.13\pm 0.02$ $1.18\pm 0.03$ $2.39\pm 0.03$
$17.52\pm 0.22$ $0.39\pm 0.01$ $0.88\pm 0.17$ $1.96\pm 0.05$
SW $9.75\pm 0.07$ $3.62\pm 0.05$ $1.19\pm 0.06$ $2.16\pm 0.01$
NE $7.44\pm 0.04$ $5.07\pm 0.05$ $0.76\pm 0.03$ $1.80\pm 0.01$
$21.40\pm 0.14$ $0.75\pm 0.01$ $2.01\pm 0.04$ $3.70\pm 0.14$
[…]
Figure 7: Radial profiles of surface brightness and temperature in the four sectorial regions marked as NE, NW, SE and SW. The surface brightness were extracted from Fig. 1. Red curves indicates the best-fit broken power-law model. Temperature increases and decreases at density discontinuities are marked by black and green vertical lines, respectively.
5 Discussion
The imaging and imaging-spectral analyses reveal detailed structures at the center of the Virgo cluster. In addition to the well-known X-ray arms, cavities, and cold fronts, a large-scale sloshing spiral pattern is identified in the GGM and residual images, connecting the two previously recognized cold fronts. Thanks to its low particle background and wide FoV, EP-FXT observations provide a continuous view of the spiral structure in the outskirts of M87, consistent with a similar pattern seen in the mosaic image from XMM-Newton (Simionescu et al. 2010). Building on these observations, we employed EP-FXT data to construct two-dimensional thermodynamic maps of the outskirts of M87, providing further insight into the large-scale physical properties of the spiral structure. In this section, we discuss the properties of the spiral structure and compare the results with those obtained by XMM-Newton.
5.1 Comparison with XMM-Newton
Figure 8: Comparison of temperature in each pixel between EP-FXT, XMM-Newton. The blue dashed line represents 1:1. The red line represents the best-fit linear regression, shown in the legend.
Using the same contour-binning regions as defined for the EP-FXT data, we performed spectral fitting on the archival XMM-Newton observation (ObsID 0803670601) over the 0.5–7.0 keV energy band, adopting the same model described in Sect. 4. The temperatures for each region are compared with those obtained from EP-FXT, as shown in Fig. 8 and Table LABEL:tab:cmp. In each panel, the blue dashed line denotes the one-to-one correspondence, while the red solid line represents the best-fit linear regression.
As shown in Fig. 8, the temperatures measured by FXT-A and FXT-B are consistent within the uncertainties, although the temperature from FXT-A appears slightly higher than that from FXT-B. This discrepancy may be attributed to the greater effective area of FXT-B at lower energies. Overall, the temperatures obtained from EP-FXT generally agree with those derived from XMM-Newton, despite minor discrepancies. These differences may result from the presence of multi-temperature components within the ICM and calibration differences of the two telescopes, which together can lead to systematic discrepancies in temperature measurements (Migkas et al. 2024; Liu et al. 2023; Schellenberger et al. 2015). Overall, EP-FXT delivers results that are consistent and comparable to those of XMM-Newton.
Table 4: Comparison of temperatures measured by EP-FXT and XMM-Newton.
Instruments X-Y Energy band (keV) a b
FXTA-FXTB 0.5-7.0 $0.85\pm 0.03$ $0.28\pm 0.04$
FXTA-XMM 0.5-7.0 $1.11\pm 0.07$ $-0.18\pm 0.13$
 [^7] Comparison of temperature fits for different instruments X and Y. Here, a and b are the linear fitting parameters.
5.2 Giant sloshing spiral
Based on the above analysis, we further confirmed the previously identified spiral structure using the GGM and residual images from EP-FXT, which provide a comprehensive view of the feature. Subsequently, we performed imaging-spectral analysis to construct two-dimensional maps of temperature, metallicity, and entropy on the outskirts of M87, providing a spatially resolved characterization of the thermodynamic properties of the spiral structure. The results show that along the extent of the spiral, metallicity is enhanced, while temperature and entropy are reduced. In addition, the spiral structure exhibits significant surface brightness enhancements, with cooler gas preferentially concentrated in the brighter regions, forming an alternating pattern of bright, cold areas and dimmer, hotter regions. Numerical simulations by Roediger et al. (2011) predict that such structures should exhibit low temperatures, high densities, and elevated metallicity. Our observations are consistent with these predictions, supporting the scenario in which the large-scale spiral structure is produced by gas sloshing triggered by a minor merger event (Ascasibar & Markevitch 2006).
Figure 9 shows the large spiral structure in Virgo, detected by EP-FXT, consisting of an inner ring-like structure and outer spiral arms, which are delineated by a white spiral line. In the northeast of the spiral structure (as indicated by the white arrow), the X-ray surface brightness is lower than in the surrounding regions, forming a concave feature. Two additional X-ray concave regions are also present in the west-southwest. This structure is similar to the Kelvin-Helmholtz rolls presented in the simulations and the smeared-out spiral structures observed in A496 (Walker et al. 2017; Ghizzardi et al. 2014). This similarity may be attributed to the influence of the Kelvin-Helmholtz instability (KHI), which generates turbulence and disturbances near the cold front, potentially affecting the formation of the spiral structure.
Overall, EP-FXT provides a clear and comprehensive view of the large-scale spiral pattern in M87. The spiral feature appears to rotate counterclockwise, extending from the cold front in the southeast to the cold front in the northwest, and further spiraling toward the cold gas clump in the southern outskirts. This spiral feature is associated with low temperature, low entropy, and high metallicity, consistent with expectations for gas sloshing phenomena. The prominent and well-defined morphology revealed by EP-FXT offers valuable insights into the formation and evolution of spiral structures in galaxy clusters.
6 Conclusion
This paper provides a detailed analysis of the images and spectra of the Virgo cluster with EP-FXT. In the images, we present a prominent and complete large-scale spiral structure. Furthermore, we produce two-dimensional maps of temperature, entropy, and metallicity that clearly illustrate the characteristics of the M87 spiral structure. The main results of this work are summarized as follows:
  1. 1.
    In the GGM image, multiple structures are identified. Two prominent X-ray arms are clearly visible: the outer side of the southwest arm bends toward the southeast, while the east arm bends toward the northwest. A shock from the northwest appears to connect to the east arm. Furthermore, cold fronts in the southeast and northwest are also readily discernible. These structures are broadly consistent with previous findings.
  2. 2.
    Based on the residual map (see Fig. 4), an ring-like structure with enhanced surface brightness is found in the outer X-ray arms. This ring-like structure, connected to the northwestern cold front, rotates counterclockwise, forming a large, outward-extending spiral structure. This structure exhibits low temperature, low entropy, and high density, which are characteristics consistent with gas sloshing induced by the merger of a gasless subcluster.
  3. 3.
    Temperatures measured with EP-FXT are consistent with those obtained with XMM-Newton within errors.
    Figure 9: The EP-FXT surface brightness residual image in the 0.3–2.5 keV band shows a spiral structure. White arrowheads indicate locations where this structure is truncated, possibly because of the KHI. The central ring structure of M87 and the outer spiral structure appear to be part of a single, larger structure, as delineated by the white spiral line.
Acknowledgements.
EP is a space mission supported by Strategic Priority Program on Space Science of Chinese Academy of Sciences, in collaboration with ESA, MPE and CNES (Grant No.XDA15310303, No.XDA15310103, No.XDA15052100). This work is supported by the International Partnership Program of Chinese Academy of Sciences,Grant No. 013GJHZ2024015FN. CJ acknowledges the National Natural Science Foundation of China through grant 12473016, and the support by the Strategic Priority Research Program of the Chinese Academy of Sciences (Grant No. XDB0550200). A.L. acknowledges the supports from the National Natural Science Foundation of China (Grant No. 12588202). A.L. is supported by the China Manned Space Program with grant no. CMS-CSST-2025-A04.

### v3

@@
 Taking advantage of the large field of view (FoV) and low particle background of EP-FXT, the image reveals a giant spiral feature connecting the cold fronts in the northwest and southeast, forming a coherent structure consistent with previous XMM-Newton and Suzaku findings.
-Furthermore, we present two-dimensional maps of temperature, metallicity, and entropy in the Virgo Cluster, covering a FoV of approximately $28.5^{\prime}$.
-These maps clearly depict a spiral structure characterized by high density, low temperature, high metallicity, and low entropy.
-These results support the scenario in which the spiral morphology originates from gas sloshing induced by a minor merger.
-In addition, the temperatures measured with EP-FXT show reasonable consistency with those obtained from XMM-Newton, within uncertainties.
+Furthermore, we present two-dimensional maps of the temperature, metallicity, and entropy in the Virgo cluster that cover a FoV of approximately $28.5^{\prime}$.
+These maps clearly depict a spiral structure that is characterized by high density, low temperature, high metallicity, and low entropy.
+These results support the scenario in which the spiral morphology originates from gas sloshing that is induced by a minor merger.
+In addition, the temperatures measured with EP-FXT are reasonably consistent with those obtained from XMM-Newton within the uncertainties.
 Key Words.:
@@
 1 Introduction
-Galaxy clusters, composed of hundreds to thousands of galaxies bound by gravity, are the largest building blocks of the cosmic web (Kravtsov & Borgani 2012; Vikhlinin et al.
+Galaxy clusters are composed of hundreds to thousands of galaxies that are bound by gravity.
+They are the largest building blocks of the cosmic web (Kravtsov & Borgani 2012; Vikhlinin et al.
 2014).
-The X-ray radiation from clusters predominantly arises from the intracluster medium (ICM), a hot, diffuse plasma where a wealth of complex structures can be seen, such as shocks, bubbles, and cold fronts, which trace the past and ongoing violent dynamic processes in the system (Markevitch & Vikhlinin 2007).
-Cold fronts, distinguished from shock fronts, are sharp interfaces separating brighter, denser, and cooler gases from the other side with a continuous pressure across the interface (Owers et al.
+The X-ray radiation from clusters predominantly arises from the intracluster medium (ICM), which is a hot, diffuse plasma with a wealth of complex structures, such as shocks, bubbles, and cold fronts that trace past and ongoing violent dynamic processes in the system (Markevitch & Vikhlinin 2007).
+Cold fronts are distinguished from shock fronts.
+They are sharp interfaces that separate brighter, denser, and cooler gases from the other side with a continuous pressure across the interface (Owers et al.
 2009; de Plaa et al.
@@
 Cold fronts or even spectacular spiral structures have been widely identified in nearby clusters (Keshet 2012).
-It is proposed that the cold front originates from a displacement of cold gas at the cluster core, triggered by gravitational perturbations—such as those induced by minor mergers (Ascasibar & Markevitch 2006; ZuHone et al.
+It is proposed that the cold front originates in a displacement of cold gas at the cluster core, triggered by gravitational perturbations, such as those induced by minor mergers (Ascasibar & Markevitch 2006; ZuHone et al.
 2010; Rossetti 2013; Markevitch et al.
@@
 2018).
-These sloshing motions can evolve into intricate spiral structures, visible in X-ray images as concentric edges (Ascasibar & Markevitch 2006; Sonkamble et al.
-2024), such as the case of A795 (Kadam et al.
+These sloshing motions can evolve into intricate spiral structures that are visible in X-ray images as concentric edges (Ascasibar & Markevitch 2006; Sonkamble et al.
+2024), for instance, A795 (Kadam et al.
 2024a).
 Alternative explanations for the sloshing-induced spiral include two-phase gas flows with differing velocities (Keshet 2012) and the rotational motion of an in-falling gas flow around a moving brightest cluster galaxy (BCG) due to the Coriolis force (Inoue 2022).
-Despite the detection of a variety of cold fronts, those manifesting themselves as a clear and complete spiral structure are rare, mainly due to difficulties in revealing large-scale and low-contrast structures with current X-ray telescopes.
+Despite the detection of a variety of cold fronts, those that manifest themselves as a clear and complete spiral structure are rare, mainly because current X-ray telescopes cannot easily reveal large-scale and low-contrast structures.
 So far, remarkable examples include those seen in A496, A795, A1775, A2029, A2052, A2566, A2657, Centaurus, NGC 7618/UGC 12491, Perseus, and Virgo (Walker et al.
@@
 2021; Kadam et al.
-2024b), mainly thanks to mosaic observations with XMM-Newton and Chandra.
-Detection and detailed investigations of large-scale sloshing spirals are essential for advancing our understanding of the physical mechanism of sloshing and dynamical evolution of galaxy clusters.
-Because of its proximity, the Virgo cluster serves as an excellent target for in-depth studies on this topic.
+2024b), which were mainly found in mosaic observations with XMM-Newton [^1] The X-ray Multi-Mirror Mission and Chandra [^2] Chandra X-ray Observatory.
+The detection and detailed investigations of large-scale sloshing spirals are essential for advancing our understanding of the physical mechanism of sloshing and of the dynamical evolution of galaxy clusters.
+Because of its proximity, the Virgo cluster is an excellent target for in-depth studies on this topic.
 Within the Virgo cluster, M87 stands out as the largest and brightest elliptical galaxy, with a redshift of 0.0042.
-It has been extensively studied across multiple wavelengths, providing valuable insights into the interaction between its central supermassive black hole ($3.2\times 10^{9}M_{\odot}$ within 3.5 pc of its center) and the surrounding ICM (Harms et al.
+It has been extensively studied at multiple wavelengths, which provided valuable insights into the interaction between its central supermassive black hole ($3.2\times 10^{9}M_{\odot}$ within 3.5 pc of its center) and the surrounding ICM (Harms et al.
 1994; Simionescu et al.
 2010; Bahcall & Sarazin 1977).
-Early X-ray observations with ROSAT, XMM-Newton, and Chandra have provided definitive evidence for active galactic nucleus (AGN) feedback surrounding the central galaxy M87 (Böhringer et al.
+Early X-ray observations with ROSAT [^3] ROentgen SATellite, XMM-Newton, and Chandra have provided definitive evidence that the feedback of an active galactic nucleus (AGN) surrounds the central galaxy M87 (Böhringer et al.
 2001; Molendi & Gastaldello 2001; Bohringer et al.
@@
 X-ray observations have revealed multiple sloshing cold fronts in the Virgo cluster.
-XMM-Newton identified cold fronts at 33 kpc and 90 kpc, and revealed a spiral structure extending to approximately 200 kpc (Simionescu et al.
+XMM-Newton identified cold fronts at 33 kpc and 90 kpc and revealed a spiral structure that extends to approximately 200 kpc (Simionescu et al.
 2010; Roediger et al.
 2011).
-Suzaku later detected additional outer cold fronts at 233 and 280 kpc and carried out detailed radial analyses of their thermodynamic properties, further confirming that gas sloshing extends well beyond the core of the cluster (Simionescu et al.
+Suzaku444originally named Astro-E2 later detected additional outer cold fronts at 233 and 280 kpc and carried out detailed radial analyses of their thermodynamic properties, which further confirmed that gas sloshing extends well beyond the core of the cluster (Simionescu et al.
 2017).
@@
 (2022), the temperature, entropy, and metallicity distributions in the outer spiral arm remain poorly characterized.
-While eROSITA is capable of covering the extended spiral structure on the outskirts of Virgo (McCall et al.
-2024), its relatively short exposure time has limited detailed analysis of the associated physical properties.
+While eROSITA [^5] The extended Roentgen Survey and Imaging Telescope Array is capable of covering the extended spiral structure in the outskirts of Virgo (McCall et al.
+2024), its relatively short exposure time has limited a detailed analysis of the associated physical properties.
 To more comprehensively investigate the role of gas sloshing and cold fronts in the outskirts of galaxy clusters, it is important to conduct observations with an X-ray telescope that offers a larger FoV and sufficient sensitivity.
 Benefiting from its large FoV and low particle background, the Follow-up X-ray Telescope on board the Einstein Probe (EP-FXT)(Yuan et al.
-2022, 2025) is well suited for detecting large-scale, low surface brightness extended structures.
-In this paper, we present deep EP-FXT observations of the Virgo cluster.
+2022, 2025) is well suited for detecting large-scale extended structures with a low surface brightness.
+We present deep EP-FXT observations of the Virgo cluster.
 Our data support the detection of the giant sloshing spiral and enable spatially resolved analyses of its thermodynamic properties.
@@
 We describe the observations in Sect.
-2, with imaging results presented in Sect.
-3 and spatially-resolved spectral results in Sect.
+2, and we present the imaging results in Sect.
+3 and the spatially resolved spectral results in Sect.
 4.
@@
 2 Observations and data reduction
-EP-FXT is one of the two payloads onboard the EP mission,which is an X-ray telescope as a result of China-Europe collaboration, launched on January 9, 2024.
-EP-FXT is capable of conducting detailed observations of targets triggered by the Wide-field X-ray Telescope (EP-WXT), as well as preselected targets designated from the ground.
-EP-FXT comprises two identical modules, FXT-A and FXT-B, each equipped with Wolter-I type X-ray optics featuring a focal length of 1.6 m and 54 nested gold-coated nickel mirror shells.
-These modules are paired with PNCCD detectors as focal-plane instruments, enabling precise measurements of photon energy, position, and timing (Cui et al.
+EP-FXT is one of the two payloads on board the EP mission, which is an X-ray telescope that is a result of China-Europe collaboration and was launched on January 9, 2024.
+EP-FXT is capable of conducting detailed observations of targets triggered by the Wide-field X-ray Telescope (EP-WXT) and of preselected targets designated from the ground.
+EP-FXT comprises two identical modules, FXT-A and FXT-B, that are each equipped with Wolter-I type X-ray optics featuring a focal length of 1.6 m and 54 nested gold-coated nickel mirror shells.
+These modules are paired with PNCCD detectors as focal-plane instruments and enable precise measurements of the photon energy, position, and timing (Cui et al.
 2023).
@@
 EP-FXT covers an energy range of 0.3–10 keV, with a FoV of $1^{\circ}\times 1^{\circ}$, a spatial resolution of $22^{\prime\prime}$ in terms of half-power diameter (HPD), and an effective area of approximately 300 $cm^{2}$ at 1 keV (one unit).
-In addition, the particle background of EP-FXT is about one-fifth that of eROSITA, and its low background characteristics make it particularly well-suited for observing low-brightness diffuse emissions (Zhang et al.
+In addition, the particle background of EP-FXT is about one-fifth of that of eROSITA, and its low background characteristics make it particularly well suited for observing low-brightness diffuse emissions (Zhang et al.
 2022).
 As the first-light object and a calibration source for EP-FXT, the Virgo cluster and its central galaxy M87 have been extensively observed.
-Due to the large FoV of EP-FXT, a single observation can effectively cover the outskirts of M87.
-All observations were performed in full-frame (FF) mode, using either the thin or medium filter.
-Three of the FXT-B observations were conducted in a closed mode, rendering the corresponding data unusable.
-As a result, only FXT-A data were used, resulting in a total effective exposure time of 295 ks.
-(see an observational log in Table LABEL:label1).
+The large FoV of EP-FXT means that a single observation can effectively cover the outskirts of M87.
+All observations were performed in full-frame (FF) mode using either the thin or medium filter.
+Three of the FXT-B observations were conducted in closed mode, which rendered the corresponding data unusable.
+As a result, only FXT-A data were used.
+This resulted in a total effective exposure time of 295 ks (see the observational log in Table LABEL:label1).
 Figure 1: Combined EP-FXT X-ray image of the Virgo cluster in the 0.3-2.5 keV band.
@@
 The bar has a length of 10.
-The arrows point to north and west.
+The arrows point north and west.
 We used the latest version 1.10 of the Follow-up X-ray Telescope Data Analysis Software (FXTDAS) [^6] http://epfxt.ihep.ac.cn/analysis developed at the EP-FXT Science Center for data reduction (Zhao et al.
@@
 New event files were generated through the fxtchain pipeline, and the out-of-time events were identified using fxtootest, which were then removed from the data with ftimgcalc [^7] https://heasarc.gsfc.nasa.gov/docs/software/lheasoft/help/heatools.html.
-When high-energy cosmic rays pass through space telescopes, they release energy onto the detectors, exciting fluorescence lines that are not modulated by the vignetting effect.
+When high-energy cosmic rays pass through space telescopes, they release energy onto the detectors, which excites fluorescence lines that are not modulated by the vignetting effect.
 To correct for this background contribution, fxtbkggen was used to generate particle backgrounds for each observation.
-We removed the particle background from each image and then used the reproject_image [^3] Derived from the Chandra Interactive Analysis of Observations (CIAO) version 4.16.0, available at https://cxc.cfa.harvard.edu/ciao/ tool to merge all exposure images along with their corresponding particle-background-subtracted images.
+We removed the particle background from each image and then used the reproject_image [^8] Derived from the Chandra Interactive Analysis of Observations (CIAO) version 4.16.0, available at https://cxc.cfa.harvard.edu/ciao/ tool to merge all exposure images along with their corresponding particle background-subtracted images.
 Subsequently, we used the merged exposure maps to correct for the vignetting effect in the particle background-subtracted X-ray images.
-Following this, the data from ObsID 13600005125, with the particle background already subtracted, were selected as the reference for the diffuse background.
-Given that the primary focus of this study is M87, we selected a background region in close proximity to the M87.
-This approach reduces the potential to underestimate the diffuse background component surrounding M87, thereby facilitating more precise background subtraction and strengthening the overall robustness of the analysis.
+Following this, the data from ObsID 13600005125, without particle background, were selected as the reference for the diffuse background.
+Because the primary focus of this study is M87, we selected a background region in close proximity to M87.
+This approach means that the diffuse background component surrounding M87 is probably not underestimated, thereby facilitating preciser background subtraction and strengthening the overall robustness of the analysis.
 We combined multiple observations, subtracted the diffuse background, and normalized the image by the maximum exposure time to produce a calibrated X-ray image.
-We used wavedetect3 with scales = ”4 8 16 32” and the default sigthresh to detect and remove point sources.
-The voids were filled using dmfilth3 with the POISSON method and a random seed of 0 to ensure reproducibility.
+We used wavedetect8 with scales = ”4 8 16 32” and the default sigthresh to detect and remove point sources.
+The voids were filled using dmfilth8 with the POISSON method and a random seed of zero to ensure reproducibility.
 Adaptive smoothing was then applied using dmimgadapt with Gaussian kernels, smoothing scales from 0.1 to 10 pixels, and a minimum of 30 counts per kernel.
-The final X-ray image in the energy range of 0.3-2.5 keV shown in Fig.
+The final X-ray image in the energy range of 0.3-2.5 keV is shown in Fig.
 1.
-Furthermore, to compare with previous results, we reanalyzed a Chandra observation (ObsID 5826) with an exposure of 126 ks, and an XMM-Newton observation (ObsID 0803670601, EPIC-MOS2) with an exposure of 65 ks.
-The data were reprocessed using the XMM-Newton Science Analysis System (SAS  [^4] https://www.cosmos.esa.int/web/xmm-newton/sas), and time intervals affected by flares were removed through high-energy light curve screening.
-In the subsequent XMM-Newton analysis, blank-sky background files were employed to improve the accuracy of background estimation and spectral analysis.
+Furthermore, to compare with previous results, we reanalyzed a Chandra observation (ObsID 5826) with an exposure of 126 ks and an XMM-Newton observation (ObsID 0803670601, EPIC-MOS2) with an exposure of 65 ks.
+The data were reprocessed using the XMM-Newton Science Analysis System (SAS  [^9] https://www.cosmos.esa.int/web/xmm-newton/sas), and time intervals affected by flares were removed through high-energy light-curve screening.
+In the subsequent XMM-Newton analysis, blank-sky background files were employed to improve the accuracy of the background estimation and spectral analysis.
 Table 1: EP-FXT observations of the Virgo cluster.
@@
 3 Imaging analysis
-To visualize faint structures in the image, we applied two different techniques for image enhancement.
-Figure 2: Adaptive GGM filtered image in the central Virgo cluster in 0.3-2.5 keV with SNR = 70.
-Prominent features are labeled.
-The two dashed line to mark the giant spiral are identical to those in Fig.
-4.
+To better reveal structures in the image, we applied multiple image enhancement techniques to enhance the visibility of both sharp and diffuse edges, thereby highlighting structural details more effectively across different spatial scales.
 Gradient filtering enhances the visibility of both sharp and flat image features, thereby facilitating their detection and analysis (Sanders et al.
 2016b, a, 2022).
-Because of the high noise in X-ray image gradient measurements within low-brightness regions, the adaptive Gaussian Gradient Magnitude (GGM) filtering is employed for dynamic scale selection.
+Because of the high noise in X-ray image gradient measurements within low-brightness regions, the adaptive Gaussian gradient magnitude (GGM) filtering was employed to select the dynamic scale.
 This method dynamically adjusts filter parameters based on local gradients, enabling better capture of image textures and edges (Sanders et al.
 2022).
-Additionally, adjusting the signal-to-noise ratio (SNR) helps smooth the image, enhances the SNR, and reveals finer details.
-Using these techniques, we applied the adaptive GGM filter to the vignetting-corrected and background-subtracted image in the 0.3–2.5 keV energy range, see Fig.
-2.
-The image clearly shows two X-ray arms, with the southwestern arm exhibiting a clockwise rotation toward the southeast (Forman et al.
+Additionally, adjusting the signal-to-noise ratio (S/N) helps smooth the image, enhances the S/N, and reveals finer details.
+Using these techniques, we applied the adaptive GGM filter to the vignetting-corrected and background-subtracted image in the 0.3–2.5 keV energy range (see Fig.
+2).
+Figure 2: Adaptive GGM filtered image in the central Virgo cluster in 0.3-2.5 keV with S/N = 70.
+Prominent features are labeled.
+The two dashed lines that mark the giant spiral are identical to those in Fig.
+4.
+The image clearly shows two X-ray arms.
+The southwestern arm exhibits a clockwise rotation toward the southeast (Forman et al.
 2007).
@@
 These features are consistent with the analysis by Sanders et al.
-(2016a) using the Chandra data.
-We also identified a region of significantly reduced surface brightness outward from the southwestern X-ray arm, which may indicate the presence of a cavity.
-The outskirts of M87 exhibit a prominent spiral morphology, extending from the northwest to the southeast, with similar spiral features also visible in the XMM-Newton mosaic image (Roediger et al.
+(2016a) based on Chandra data.
+We also identified a region in which the surface brightness was significantly reduced outward of the southwestern X-ray arm, which may indicate a cavity.
+The outskirts of M87 exhibit a prominent spiral morphology that extends from the northwest to the southeast, and similar spiral features are also visible in the XMM-Newton mosaic image (Roediger et al.
 2011; Simionescu et al.
 2010).
-We then generate a residual image to better visualize this large-scale faint feature.
-First, we need to estimate the first-order radial profile using the well-known double $\beta$ model and subtract it.
-As shown in Fig.
-3, the red solid line represents the best-fit double $\beta$ model to the EP-FXT data (black points).
-The blue points represent the XMM-Newton data and the green points represent the Chandra data, for comparison.
-As shown, apart from the central region affected by the AGN, the EP-FXT data are in agreement with the double $\beta$ model and exhibit consistency with the radial profiles measured by the three telescopes.
-Furthermore, thanks to EP-FXT’s large FoV, its observational range extends to regions up to $30^{\prime}$ from the outskirts of M87.
+We then generated a residual image to better visualize this large-scale faint feature.
+First, we estimated the first-order radial profile using the well-known double $\beta$ model and subtracted it (see Fig.
+3).
+Except for the central region that is affected by the AGN, the EP-FXT data agree with the double $\beta$ model and are consistent with the radial profiles measured by the three telescopes.
+Furthermore, the EP-FXT’s large FoV means that its observational range extends to regions up to $30^{\prime}$ from the outskirts of M87.
 Figure 3: Normalized surface brightness profiles measured with EP-FXT, XMM-Newton, and Chandra.
-The EP-FXT data are fitted with a double $\beta$ model, with individual components displayed in dashed yellow curves.
+The EP-FXT data were fit with a double $\beta$ model.
+Individual components are displayed as dashed yellow curves.
 Figure 4: Residual image obtained by subtracting the image in Fig.
 1 with the elliptical $\beta$ model.
-A Gaussian smooth with a kernel size of 7 pix is applied (1 pixel = $9.6^{\prime\prime}$).
+A Gaussian smoothing with a kernel size of 7 pix was applied (1 pixel = $9.6^{\prime\prime}$).
 Four regions of interest (NW, SE, SW and NE) are marked.
-The white dashed lines mark the position of radial discontinuity.
-The blue dashed lines mark the positions of the giant spiral arm.
+The dashed white lines mark the position of the radial discontinuity.
+The dashed blue lines mark the position of the giant spiral arm.
 To characterize the radial profile, we used Sherpa (version 4.16.0 within CIAO) [^10] https://cxc.cfa.harvard.edu/sherpa/ to model the surface brightness.
-To account for the apparent elongated morphology as seen from the image, we fitted an elliptical $\beta$ model to the data; the best‑fit parameters are listed in Table 2.
+To account for the apparently elongated morphology in the image, we fit an elliptical $\beta$ model to the data; the best‑fit parameters are listed in Table 2.
 We then subtracted this model from the original image to obtain the residual image shown in Fig.
@@
 $90.58^{+0.39}_{-0.39}$ $10^{+0.04}_{-0.04}$ $248.13^{+0.12}_{-0.12}$ $4528.09^{+17.9}_{-17.9}$ $74^{+0.09}_{-0.09}$
-The residual image reveals a few patterns, including X-ray arms and discontinuities which indicate cold fronts or shocks.
-Notably, we identified a ring-like structure outside the X-ray arms, which displays a break in the southern part.
+The residual image revealed a few patterns, including X-ray arms and discontinuities, which indicate cold fronts or shocks.
+Notably, we identified a ring-like structure outside the X-ray arms that displays a break in the southern part.
 This may be attributed to interactions between the AGN jets and the ICM (Forman et al.
 2005).
-Most importantly, the large-scale spiral structure with excessive brightness is evident on the outskirts, spiraling from the cold front in the northwest to the southern cold gas clump, marked with blue dashed lines in Fig.
-4 and white dashed lines in Fig.
-2.
+Most importantly, a large-scale spiral structure with an excessive brightness is evident in the outskirts, where it spirals from the cold front in the northwest to the southern cold gas clump (marked with dashed blue lines in Fig.
+4 and dashed white lines in Fig.
+2).
 Consistent with the mosaic XMM-Newton results (Simionescu et al.
-2010), the EP-FXT observations clearly reveal the spiral features as continuous and well-defined across the entire FoV, likely owing to the instrument’s low particle background and wide spatial coverage.
-The northwestern outer edge (a cold front) of the spiral lies approximately $18.3^{\prime}$ from the center.
-The structure rotates counterclockwise toward the south, with its southern inner edge located approximately $17.5^{\prime}$ from the center.
-Figure 5: An example of the EP-FXT spectrum from a contour-binning region in M87.
+2010), the EP-FXT observations clearly reveal that the spiral features are continuous and well defined across the entire FoV, likely owing to the low particle background and wide spatial coverage of the instrument.
+The outer northwestern edge (a cold front) of the spiral lies approximately $18.3^{\prime}$ away from the center.
+The structure rotates counterclockwise toward the south, and its inner southern edge is located approximately $17.5^{\prime}$ away from the center.
+Figure 5: EP-FXT spectrum from a contour-binning region in M87.
 The black and red points represent the spectral data from FXT-A and FXT-B for ObsID 13600005124, with the corresponding background from ObsID 13600005125 subtracted, and the solid line represents the best-fit model.
 The green and blue points represent the spectra from FXT-A and FXT-B for ObsID 13600005125, respectively.
-The light cyan and orange represent the particle backgrounds for ObsIDs 13600005124 and 13600005125, respectively.
-Figure 6: Temperature, pressure, entropy, and metallicty maps measured with EP-FXT, pixelated with the contour-binning method. The white dashed curves mark the spiral structure as in Fig. 4.
-4 Imaging-spectral analysis
-The X-ray image obtained with EP-FXT was divided into 142 regions using the contour-binning software (Sanders 2006) with a SNR of 300.
-We adopted a flat-weighting assumption, assuming that the X-ray emission within each region is spatially uniform, for the Ancillary Response Files (ARFs) calculation.
+The light cyan and orange points represent the particle backgrounds for ObsIDs 13600005124 and 13600005125, respectively.
+Figure 6: Temperature, pressure, entropy, and metallicty maps measured with EP-FXT, pixeled with the contour-binning method. The dashed white curves mark the spiral structure as in Fig. 4.
+4 Imaging spectral analysis
+The X-ray image obtained with EP-FXT was divided into 142 regions using the contour-binning software (Sanders 2006) with a S/N of 300.
+We adopted a flat-weighting assumption, assuming that the X-ray emission within each region is spatially uniform, to calculate the ancillary response files (ARFs).
 The ARFs were generated using the EP-FXT response tool fxtarfgen.
-During the calculation, the parameter extend was set to 1 to account for the extended nature of the source, thus more accurately capturing the spatial distribution of the emission within each extraction region.
-The spectra were grouped using grppha with a minimum of 10 counts per bin to ensure the applicability of the C-statistic.
+During the calculation, the parameter extend was set to one to account for the extended nature of the source, thus more accurately capturing the spatial distribution of the emission within each extraction region.
+The spectra were grouped using grppha with a minimum of ten counts per bin to ensure the applicability of the C-statistic.
 To minimize the impact of background contamination on the source spectra and improve the accuracy of the spectral analysis, we selected observations of the M87 source and its corresponding background, ensuring similar observation times, the same filter settings, and longer exposure durations; specifically, ObsID 13600005124 and ObsID 13600005125.
@@
 Therefore, we selected the spectra from the same CCD region in ObsID 13600005125 to serve as the background for ObsID 13600005124.
-For regions in ObsID 13600005125 that are contaminated by bright point sources or extended sources, background data from nearby uncontaminated areas were used instead.
-Given that our analysis focuses on M87, we intentionally retained the Virgo cluster component in the background data because it more accurately reflects the diffuse emission environment surrounding the galaxy.
-Spectral fitting of the combined FXT-A and FXT-B spectra in each region was performed using XSPEC version 12.14.1 over the 0.5–7.0 keV energy range.
+For regions in ObsID 13600005125 that were contaminated by bright point sources or extended sources, background data from nearby uncontaminated areas were used instead.
+Because our analysis focuses on M87, we intentionally retained the Virgo cluster component in the background data because it reflects the diffuse emission environment surrounding the galaxy better.
+We spectrally fit the combined FXT-A and FXT-B spectra in each region using XSPEC version 12.14.1 over the 0.5–7.0 keV energy range.
 The diffuse ICM thermal emission was described with the APEC model (version 3.0.9), and the Galactic hydrogen absorption was accounted for using the TBABS model, adopting the solar abundance table from Anders & Grevesse (1989), with the redshift fixed at 0.0042 and the hydrogen column density fixed at $2.11\times 10^{20}\mathrm{cm}^{-2}$ (Willingale et al.
 2013).
-Other parameters are set to be free.
+Other parameters were set to be free.
 We examined the C-Stat/d.o.f.
 values for all regions, which were consistently around 1.2.
-We plotted the spectrum of one region, as shown in Fig.
+We plot the spectrum of one region in Fig.
 5.
-The region belongs to the contour-binning partition and is located in a region approximately $6.6^{\prime}$ from the cluster center.
+The region belongs to the contour-binning partition and is located in a region approximately $6.6^{\prime}$ away from the cluster center.
 The red and black points represent the spectral data from FXT‐A and FXT‐B in the source region (with the diffuse background subtracted), while the solid line denotes the best‐fit model.
@@
 6) reveals a central region in M87 with temperatures below 2.5 keV, along with two distinct low-temperature regions in the northwest and southeast.
-Sharp temperature increases are seen exterior to these regions, coincident with the locations of two concentric, arc-shaped cold fronts.
-A low-temperature belt extends outwards from the central region, connecting the low-temperature regions in the southeast and northwest, and spiraling counterclockwise from the northwest, forming a large-scale low-temperature spiral structure.
-In addition, a region with a lower temperature, compared to the surrounding areas, is located on the outer side of the southwest arm.
+Sharp temperature increases are seen outside these regions that are coincident with the locations of two concentric arc-shaped cold fronts.
+A low-temperature belt extends outward from the central region.
+It connects the low-temperature regions in the southeast and northwest and spirals in counterclockwise from the northwest, forming a large-scale low-temperature spiral structure.
+In addition, a region with a lower temperature than the surrounding areas is located outside of the southwest arm.
 A detailed analysis is presented in Sect.
 4.2.
-The metallicity are obtained from the APEC model.
-The metallicity is lower within the two X-ray arms but higher to the north and south of the center, which is consistent with the findings in Simionescu et al.
+The metallicity was obtained from the APEC model.
+The metallicity is lower within the two X-ray arms, but higher north and south of the center, which is consistent with the findings by Simionescu et al.
 (2007); Gatuzz et al.
@@
 Additionally, beyond $12^{\prime}$, the metallicity in the northwest is significantly higher than that in the southeast, likely due to the accumulation of cold gas in the northwest.
-In addition, the metallicity of the spiral structure displays higher values, further supporting the existence of the spiral structure.
-To investigate its thermodynamic properties, we examine the distributions of pseudo-pressure and pseudo-entropy.
+In addition, the metallicity of the spiral structure displays higher values, which further supports the existence of the spiral structure.
+To investigate its thermodynamic properties, we examined the distributions of pseudo-pressure and pseudo-entropy.
 The thermal pressure and entropy can be defined (Sasaki et al.
-2016) as $P=kTn_{\textnormal{e}}$ and $S=kTn_{\textnormal{e}}^{-2/3}$, respectively, where $n_{e}$ is the electron density and $kT$ is the temperature.
-Based on the emission measure obtained from spectral fits, we obtained pseudo-entropy and pseudo-pressure.
-we convert $P$ and $S$ to $P\propto kT{\rm norm}^{0.5}$ and $S\propto kT{\rm norm}^{-1/3}$, respectively (Fabian et al.
+2016) as $P=kTn_{\textnormal{e}}$ and $S=kTn_{\textnormal{e}}^{-2/3}$, respectively, where $n_{e}$ is the electron density, and $kT$ is the temperature.
+Based on the emission measure obtained from spectral fits, we obtained the pseudo-entropy and pseudo-pressure.
+We convert $P$ and $S$ into $P\propto kT{\rm norm}^{0.5}$ and $S\propto kT{\rm norm}^{-1/3}$, respectively (Fabian et al.
 2006).
@@
 6.
-In the central region of M87, the pressure map reveals decreasing pressure along the eastern and southwestern X-ray arms, and increasing pressure in the northwest and southeast.
-The entropy map shows low-entropy features coincident with both X-ray arms, exhibiting an entropy asymmetry towards the northwest and southeast near the core.
-These results are consistent with those reported in Simionescu et al.
+In the central region of M87, the pressure map reveals decreasing pressure along the eastern and southwestern X-ray arms and increasing pressure in the northwest and southeast.
+The entropy map shows low-entropy features coincident with both X-ray arms, exhibiting an entropy asymmetry northwest and southeast near the core.
+These results are consistent with those reported by Simionescu et al.
 (2007).
-The outer spiral structure also displays lower entropy values, further supporting the existence of the spiral structure.
-To sum up, the EP-FXT observations revealed the spiral structure with high completeness and clarity.
-Moreover, we systematically mapped its thermodynamic properties—including temperature, metal abundance, and entropy—through two-dimensional distributions.
+The outer spiral structure also displays lower entropy values, which further supports the existence of the spiral structure.
+To summarize, the EP-FXT observations revealed the spiral structure with high completeness and clarity.
+Moreover, we systematically mapped its thermodynamic properties, including temperature, metal abundance, and entropy, through two-dimensional distributions.
 These results provide detailed and spatially resolved evidence for the existence and evolution of the spiral structure.
-4.2 Sectorial radial profiles
+4.2 Radial sector profiles
 In the residual map (see Fig.
-4), we defined four sectorial regions of interest (NW, SE, NE and SW).
+4), we defined four sector regions of interest (NW, SE, NE, and SW).
 Using Fig.
-1, radial profiles of the surface brightness and temperture are extracted in each region (see Fig.
+1, we extracted the radial profiles of the surface brightness and temperature in each region (see Fig.
 7).
-The surface brightness profiles near the cold fronts are fitted with a broken power-law.
+The surface brightness profiles near the cold fronts were fit with a broken power law.
 The best‐fit parameters of the broken power‐law model are listed in Table 3.
-In the northeast, two surface brightness drops are detected at $R=7.44\pm 0.04^{\prime}$ and $R=21.40\pm 0.14^{\prime}$, respectively.
-At $R=21.40\pm 0.14^{\prime}$, a modest increase in temperature is found from $2.16^{+0.11}_{-0.10}\leavevmode\nobreak\ \mathrm{keV}$ to $2.50^{+0.19}_{-0.16}\leavevmode\nobreak\ \mathrm{keV}$ at the $1.7\sigma$ level (The uncertainties are quoted at the $1\sigma$ level).
-While the temperature change alone is not highly significant, the corresponding surface brightness discontinuity and the location in the outskirts of M87—where the SNR is lower—suggest that this feature likely marks the outer boundary of the northeastern spiral structure.
-In the northwest, a drop in surface brightness is found at $R=18.33\pm 0.08^{\prime}$, accompanied by a rise in temperature from $2.39^{+0.06}_{-0.07}$ to $2.70^{+0.12}_{-0.14}\leavevmode\nobreak\ \mathrm{keV}$, corresponding to a significance of approximately $2.2\sigma$.
-Despite the moderate significance of the temperature jump, the presence of a surface brightness edge, coupled with the low SNR in the outskirts, supports the classification of this feature as a cold front, in agreement with previous findings.
-In the southeast, a cold front is identified at $R=6.76\pm 0.03^{\prime}$, indicated by a drop in surface brightness and a temperature increase from $2.43^{+0.03}_{-0.02}$ to $2.58^{+0.04}_{-0.03}\leavevmode\nobreak\ \mathrm{keV}$ at the $3.3\sigma$ level.
+In the northeast, two surface brightness drops are detected at $R=7.44\pm 0.04^{\prime}$ and $R=21.40\pm 0.14^{\prime}$.
+At $R=21.40\pm 0.14^{\prime}$, a modest increase in the temperature is found from $2.16^{+0.11}_{-0.10}\leavevmode\nobreak\ \mathrm{keV}$ to $2.50^{+0.19}_{-0.16}\leavevmode\nobreak\ \mathrm{keV}$ at the $1.7\sigma$ level (the uncertainties are quoted at the $1\sigma$ level).
+While the temperature change alone is not highly significant, the corresponding surface brightness discontinuity and the location in the outskirts of M87 (where the S/N is lower) suggest that this feature likely marks the outer boundary of the northeastern spiral structure.
+In the northwest, a drop in the surface brightness is found at $R=18.33\pm 0.08^{\prime}$, accompanied by a rise in temperature from $2.39^{+0.06}_{-0.07}$ to $2.70^{+0.12}_{-0.14}\leavevmode\nobreak\ \mathrm{keV}$, corresponding to a significance of approximately $2.2\sigma$.
+Despite the moderate significance of the temperature jump, the surface brightness edge, coupled with the low S/N in the outskirts, supports the classification of this feature as a cold front.
+This agrees with previous findings.
+In the southeast, a cold front is identified at $R=6.76\pm 0.03^{\prime}$, indicated by a drop in the surface brightness and a temperature increase from $2.43^{+0.03}_{-0.02}$ to $2.58^{+0.04}_{-0.03}\leavevmode\nobreak\ \mathrm{keV}$ at the $3.3\sigma$ level.
 At $R=17.52\pm 0.22^{\prime}$, the surface brightness shows a mild discontinuity, transitioning from a deficit to an excess in the residual map (see Fig.
@@
 This feature may correspond to the inner boundary of the southeastern portion of the spiral structure.
-A region of excess surface brightness is detected in the southwest at $R=9.75\pm 0.07^{\prime}$, potentially associated with an arc-shaped feature from past AGN activity (Forman et al.
+A region of excess surface brightness is detected in the southwest at $R=9.75\pm 0.07^{\prime}$, which might be associated with an arc-shaped feature from past AGN activity (Forman et al.
 2005).
@@
 The two profiles show a broadly consistent trend, with comparable shapes and temperatures ranging from approximately 2 to 3 keV.
-Table 3: Best-fit broken power-law model parameters.
+Table 3: Parameters of the best-fit broken power-law model.
 region $R_{\rm break}$ (arcmin) $I_{0}$ ( $10^{-2}$ cts s-1 arcmin-2) $\alpha_{1}$ $\alpha_{2}$
@@
 […]
-Figure 7: Radial profiles of surface brightness and temperature in the four sectorial regions marked as NE, NW, SE and SW.
-The surface brightness were extracted from Fig.
+Figure 7: Radial profiles of surface brightness and temperature in the four sectorial regions marked NE, NW, SE, and SW.
+The surface brightness was extracted from Fig.
 1.
-Red curves indicates the best-fit broken power-law model.
-Temperature increases and decreases at density discontinuities are marked by black and green vertical lines, respectively.
+The red curves indicate the best-fit broken power-law model.
+The temperature increases and decreases at density discontinuities are marked by vertical black and green lines, respectively.
 5 Discussion
-The imaging and imaging-spectral analyses reveal detailed structures at the center of the Virgo cluster.
-In addition to the well-known X-ray arms, cavities, and cold fronts, a large-scale sloshing spiral pattern is identified in the GGM and residual images, connecting the two previously recognized cold fronts.
-Thanks to its low particle background and wide FoV, EP-FXT observations provide a continuous view of the spiral structure in the outskirts of M87, consistent with a similar pattern seen in the mosaic image from XMM-Newton (Simionescu et al.
+The imaging and imaging-spectral analyses revealed detailed structures at the center of the Virgo cluster.
+In addition to the well-known X-ray arms, cavities, and cold fronts, a large-scale sloshing spiral pattern is identified in the GGM and residual images, where it connects the two previously recognized cold fronts.
+The low particle background and wide FoV allow the EP-FXT observations to provide a continuous view of the spiral structure in the outskirts of M87.
+This is consistent with a similar pattern seen in the mosaic image from XMM-Newton (Simionescu et al.
 2010).
-Building on these observations, we employed EP-FXT data to construct two-dimensional thermodynamic maps of the outskirts of M87, providing further insight into the large-scale physical properties of the spiral structure.
+Building on these observations, we employed EP-FXT data to construct two-dimensional thermodynamic maps of the outskirts of M87 to provide further insight into the large-scale physical properties of the spiral structure.
 In this section, we discuss the properties of the spiral structure and compare the results with those obtained by XMM-Newton.
 5.1 Comparison with XMM-Newton
-Figure 8: Comparison of temperature in each pixel between EP-FXT, XMM-Newton. The blue dashed line represents 1:1. The red line represents the best-fit linear regression, shown in the legend.
+Figure 8: Comparison of the temperature in each pixel between EP-FXT and XMM-Newton.
+The dashed blue line represents the 1:1 ratio.
+The red line represents the best-fit linear regression, shown in the legend.
 Using the same contour-binning regions as defined for the EP-FXT data, we performed spectral fitting on the archival XMM-Newton observation (ObsID 0803670601) over the 0.5–7.0 keV energy band, adopting the same model described in Sect.
 4.
-The temperatures for each region are compared with those obtained from EP-FXT, as shown in Fig.
+The temperatures for each region were compared with those obtained from EP-FXT, as shown in Fig.
 8 and Table LABEL:tab:cmp.
@@
 As shown in Fig.
-8, the temperatures measured by FXT-A and FXT-B are consistent within the uncertainties, although the temperature from FXT-A appears slightly higher than that from FXT-B.
+8, the temperatures measured by FXT-A and FXT-B are consistent within the uncertainties, although the temperature from FXT-A appears to be slightly higher than that from FXT-B.
 This discrepancy may be attributed to the greater effective area of FXT-B at lower energies.
 Overall, the temperatures obtained from EP-FXT generally agree with those derived from XMM-Newton, despite minor discrepancies.
-These differences may result from the presence of multi-temperature components within the ICM and calibration differences of the two telescopes, which together can lead to systematic discrepancies in temperature measurements (Migkas et al.
+These differences may result from the multitemperature components within the ICM and from calibration differences of the two telescopes, which together can lead to systematic discrepancies in the temperature measurements (Migkas et al.
 2024; Liu et al.
@@
 2015).
-Overall, EP-FXT delivers results that are consistent and comparable to those of XMM-Newton.
-Table 4: Comparison of temperatures measured by EP-FXT and XMM-Newton.
+Overall, EP-FXT delivers results that are consistent with and comparable to those of XMM-Newton.
+Table 4: Comparison of the temperatures measured by EP-FXT and XMM-Newton.
 Instruments X-Y Energy band (keV) a b
@@
 Based on the above analysis, we further confirmed the previously identified spiral structure using the GGM and residual images from EP-FXT, which provide a comprehensive view of the feature.
-Subsequently, we performed imaging-spectral analysis to construct two-dimensional maps of temperature, metallicity, and entropy on the outskirts of M87, providing a spatially resolved characterization of the thermodynamic properties of the spiral structure.
-The results show that along the extent of the spiral, metallicity is enhanced, while temperature and entropy are reduced.
-In addition, the spiral structure exhibits significant surface brightness enhancements, with cooler gas preferentially concentrated in the brighter regions, forming an alternating pattern of bright, cold areas and dimmer, hotter regions.
+Subsequently, we performed an imaging-spectrum analysis to construct two-dimensional maps of the temperature, metallicity, and entropy in the outskirts of M87, providing a spatially resolved characterization of the thermodynamic properties of the spiral structure.
+The results show that throughout the spiral, the metallicity is enhanced, while temperature and entropy are reduced.
+In addition, the spiral structure exhibits significant surface brightness enhancements.
+Cooler gas is preferentially concentrated in the brighter regions, where it forms an alternating pattern of bright cold areas and dimmer hotter regions.
 Numerical simulations by Roediger et al.
-(2011) predict that such structures should exhibit low temperatures, high densities, and elevated metallicity.
-Our observations are consistent with these predictions, supporting the scenario in which the large-scale spiral structure is produced by gas sloshing triggered by a minor merger event (Ascasibar & Markevitch 2006).
-Figure 9 shows the large spiral structure in Virgo, detected by EP-FXT, consisting of an inner ring-like structure and outer spiral arms, which are delineated by a white spiral line.
-In the northeast of the spiral structure (as indicated by the white arrow), the X-ray surface brightness is lower than in the surrounding regions, forming a concave feature.
+(2011) predicted that these structures should exhibit low temperatures, high densities, and an elevated metallicity.
+Our observations are consistent with these predictions and support the scenario in which the large-scale spiral structure is produced by gas sloshing that was triggered by a minor merger event (Ascasibar & Markevitch 2006).
+Figure 9 shows the large spiral structure in Virgo as detected by EP-FXT.
+It consists of an inner ring-like structure and outer spiral arms.
+Northeast of the spiral structure (as indicated by the white arrow), the X-ray surface brightness is lower than in the surrounding regions and forms a concave feature.
 Two additional X-ray concave regions are also present in the west-southwest.
@@
 2014).
-This similarity may be attributed to the influence of the Kelvin-Helmholtz instability (KHI), which generates turbulence and disturbances near the cold front, potentially affecting the formation of the spiral structure.
+This similarity may be attributed to the influence of the Kelvin-Helmholtz instability, which generates turbulence and disturbances near the cold front that might affect the formation of the spiral structure.
 Overall, EP-FXT provides a clear and comprehensive view of the large-scale spiral pattern in M87.
-The spiral feature appears to rotate counterclockwise, extending from the cold front in the southeast to the cold front in the northwest, and further spiraling toward the cold gas clump in the southern outskirts.
-This spiral feature is associated with low temperature, low entropy, and high metallicity, consistent with expectations for gas sloshing phenomena.
+The spiral feature appears to rotate counterclockwise and extends from the cold front in the southeast to the cold front in the northwest.
+It further spirals toward the cold gas clump in the southern outskirts.
+This spiral feature is associated with a low temperature, low entropy, and high metallicity, consistent with expectations for gas-sloshing phenomena.
 The prominent and well-defined morphology revealed by EP-FXT offers valuable insights into the formation and evolution of spiral structures in galaxy clusters.
 6 Conclusion
-This paper provides a detailed analysis of the images and spectra of the Virgo cluster with EP-FXT.
-In the images, we present a prominent and complete large-scale spiral structure.
-Furthermore, we produce two-dimensional maps of temperature, entropy, and metallicity that clearly illustrate the characteristics of the M87 spiral structure.
-The main results of this work are summarized as follows:
+We provided a detailed analysis of the images and spectra of the Virgo cluster with EP-FXT.
+The images show a prominent and complete large-scale spiral structure.
+Furthermore, we produced two-dimensional maps of the temperature, entropy, and metallicity that clearly illustrated the characteristics of the M87 spiral structure.
+The main results of this work are summarized below.
   1. 1.
     In the GGM image, multiple structures are identified.
-Two prominent X-ray arms are clearly visible: the outer side of the southwest arm bends toward the southeast, while the east arm bends toward the northwest.
-A shock from the northwest appears to connect to the east arm.
+Two prominent X-ray arms are clearly visible: The outer side of the southwest arm bends toward the southeast, and the east arm bends toward the northwest.
+A shock from the northwest appears to be connected to the east arm.
 Furthermore, cold fronts in the southeast and northwest are also readily discernible.
@@
     Based on the residual map (see Fig.
-4), an ring-like structure with enhanced surface brightness is found in the outer X-ray arms.
-This ring-like structure, connected to the northwestern cold front, rotates counterclockwise, forming a large, outward-extending spiral structure.
-This structure exhibits low temperature, low entropy, and high density, which are characteristics consistent with gas sloshing induced by the merger of a gasless subcluster.
+4), a ring-like structure with an enhanced surface brightness is found in the outer X-ray arms.
+This ring-like structure is connected to the northwestern cold front and rotates counterclockwise.
+It forms a large, outward-extending spiral structure.
+This structure has a low temperature, low entropy, and high density.
+These characteristics are consistent with gas sloshing induced by the merger of a gasless subcluster.
   3. 3.
-    Temperatures measured with EP-FXT are consistent with those obtained with XMM-Newton within errors.
-    Figure 9: The EP-FXT surface brightness residual image in the 0.3–2.5 keV band shows a spiral structure.
-White arrowheads indicate locations where this structure is truncated, possibly because of the KHI.
-The central ring structure of M87 and the outer spiral structure appear to be part of a single, larger structure, as delineated by the white spiral line.
+    The temperatures measured with EP-FXT are consistent with those obtained with XMM-Newton within the errors.
+    Figure 9: EP-FXT surface brightness residual image in the 0.3–2.5 keV band showing a spiral structure.
+The white arrowheads indicate locations where this structure is truncated, possibly because of the Kelvin-Helmholtz instability.
+The central ring structure of M87 and the outer spiral structure appear to be part of a single larger structure, as delineated by the white spiral line.
 Acknowledgements.